Hello,
Can you help me solve this problem? A woman made a deposit of $261. If her deposit consisted of 77 bills, some of them one-dollar bills and the rest being five-dollar bills, how many one-dollar bills did she deposit?
Thanks,
KC

Answer by ankor@dixie-net.com(12702)

A woman made a deposit of $261. If her deposit consisted of 77 bills, some of them one-dollar bills and the rest being five-dollar bills, how many one-dollar bills did she deposit?
:
Let x = number of $1 bills; Let y = number of $5 bills
:
x + y = 77
y = (77 - x)
:
1x + 5y = 261
:
Substitute (77-x) for y in the above equation:
x + 5(77-x) = 261
:
x + 385 - 5x = 261
:
x - 5x = 261 - 385
:
-4x = -124
:
x = -124/-4
:
x = + 31 ea $1 bills
:
y = 77 - 31
:
y = 46 ea $5 bills
;
;
Check: 31 + 5(46) = $261